# Oriented Object Detection in Optical Remote Sensing Images using Deep Learning: A Survey

### v4

@@ 6.2 Large Aspect Ratio
RS images commonly contain some categories with an extremely large aspect ratio, such as bridges, ships, and harbors. The RIoU between these categories with anchors is heavily sensitive to the orientation error, thereby causing two typical challenges, i.e., spatial misalignment and inaccurate localization. Accordingly, a series of well-designed assignment strategies and regression loss are proposed to remedy corresponding issues. Zhu et al. (2020) propose length-independent IoU (LiIoU) that intercepts part of the object box along its long side based on the length of the anchor and then calculates the IoU between the intercepted object box and anchor. The LiIoU contributes to assigning more positive samples for long objects than naïve IoU, thus improving the recall rate. Following ATSS (Zhang et al., 2020b), SASM (Hou et al., 2022) design a monotonic decreasing function of aspect ratio as a weight for the IoU threshold that controls the sample assignment, enabling the long objects to be assigned a low IoU threshold. Additionally, several methods (Liu et al., 2022b; Qiao et al., 2023) construct the weighted orientation loss depending on aspect ratio, which can decrease the effect of aspect ratio on orientation regression.
6.3 Complex Backgrounds
Due to the wide visual field and complex earth’s surface, RS images typically contain a variety of complex backgrounds, causing significant interference in detection. Objects are frequently surrounded by different backgrounds, requiring detectors to possess sufficient discrimination. Besides, there may be backgrounds containing textures and shapes similar to objects, causing a large number of false alarms. To this end, a series of efforts have been made to suppress background noise and emphasize the valuable areas of the objects. CFC-Net (Ming et al., 2022) and PETDet (Li et al., 2024) both combine channel attention and spatial attention modules to learn the semantic correlation between foreground and background. Such a strategy follows a self-attention mechanism, thus lacking the guidance of the foreground regions. To enhance the discrimination of foreground, FoRDet (Zhang et al., 2022) proposes a foreground relation module for foreground-contextual representations under the supervision of the designed foreground map. Similarly, CBDA-Net (Liu et al., 2022b) builds two parallel spatial attention streams to capture center and boundary attention features, which could assist detectors in improving object localization accuracy. Besides, SCRDet (Yang et al., 2019a) adopts the pixel attention network to generate a saliency map that can separate foreground from background, and use Squeeze-and-Excitation (SE) blocks (Hu et al., 2020) as the channel attention network to enhance the saliency map further. Following this paradigm, SCRDet++ (Yang et al., 2022a) proposes an instance-level denoising (InLD) module that weakens the feature response of the background region while decouples the features of different categories into their corresponding channels.
@@ 6.4 Discussion
Scale variations, large aspect ratio, and complex background are crucial issues in RS object detection tasks, which evolve more severely when it comes to oriented object detection tasks. As stated above, a number of works have been proposed to tackle these issues from various perspectives, e.g., data augmentation, assignment strategies, re-weighted orientation loss, and attention mechanism. Unfortunately, exploration into solutions for these issues is far from mature and so further research may be beneficial. For instance, there is still a significant performance gap in detecting small/long and normal objects even for state-of-the-art detectors. On the other hand, the general split-and-detect scheme is inefficient during inference due to too many empty patches that only contain background. Several prior works have provided preliminary consideration on these points, e.g., super-resolution-based object detection (Shermeyer and Van Etten, 2019; Liu et al., 2023; Zhang et al., 2023a), focus-and-detect schemes (Duan et al., 2021; Koyun et al., 2022).
7 Datasets and Performance Evaluation
@@ 7.1 Datasets
As a data-driven technology, deep learning is inseparable from various datasets (Everingham et al., 2010; Russakovsky et al., 2015; Lin et al., 2014; Kuznetsova et al., 2020). Throughout the history of object detection, datasets have played an indispensable role not only in training models but also as an impartial platform for evaluating and verifying various methods (Liu et al., 2020). With the rapid development of earth observation technologies, an enormous amount of high-quality RS images can be easily collected to build large-scale datasets. Recently, several research groups have released dozens of high-quality RS image datasets, each of which dramatically boosts the development of RS object detection. Datasets annotated only with HBBs are not covered here, including DIOR (Li et al., 2020), LEVIR (Zou and Shi, 2018), NWPU VHR-10 (Cheng et al., 2014), RSOD (Xiao et al., 2015; Long et al., 2017), xView (Lam et al., 2018), and HRRSD (Zhang et al., 2019). In this subsection, we only focus on introducing RS datasets annotated with OBBs, including SZTAKI-INRIA (Benedek et al., 2012), 3K vehicle (Liu and Mattyus, 2015), UCAS-AOD (Zhu et al., 2015), VEDAI (Razakarivony and Jurie, 2016), HRSC2016 (Liu et al., 2016b), DOTA (Xia et al., 2018; Ding et al., 2022), ShipRSImageNet (Zhang et al., 2021b), and DIOR-R (Cheng et al., 2022a)]. Table 2 statistics the parameters of the above RS datasets for intuitive comparison, and only the most typical among them are described in detail due to the space restriction. For more details please refer to Section C of Appendix.
Dataset |
Category |
Quantity |
Instance |
GSD |
Resolution |
SZTAKI-INRIA (Benedek et al., 2012) |
1 |
9 |
665 |
3K vehicle (Liu and Mattyus, 2015) |
1 |
20 |
14235 |
0.13m |
VEDAI (Razakarivony and Jurie, 2016) |
9 |
1210 |
3640 |
0.125m |
UCAS-AOD (Zhu et al., 2015) |
2 |
2420 |
14596 |
HRSC2016 (Liu et al., 2016b) |
25 |
1070 |
2976 |
0.42m |
DOTA-V1.0 (Xia et al., 2018) |
15 |
2806 |
188282 |
0.14.5m |
DOTA-V1.5 |
16 |
2806 |
403318 |
0.14.5m |
DOTA-V2.0 (Ding et al., 2022) |
18 |
11268 |
1793658 |
0.14.5m |
FGSD (Chen et al., 2020) |
43 |
5634 |
2612 |
0.121.93m |
ShipRSImageNet (Zhang et al., 2021b) |
50 |
3435 |
17573 |
0.126m |
DIOR-R (Cheng et al., 2022a) |
20 |
23463 |
192518 |
0.530m |
@@
DOTA (Xia et al., 2018; Ding et al., 2022) contains large quantities of objects with a considerable variety of orientations, scales, and appearances. The images are selected from different sensors and platforms, including Google Earth, GF-2 Satellite, and UAVs. There are three versions of this dataset. The number of images and instances in three versions of DOTA are summarized in Table 3. DOTA-V1.0 (Xia et al., 2018) and DOTA-V1.5 share the same images, which are split into training, validation, and test subsets. As an extension of DOTA-V1.0, DOTA-V1.5 annotates extremely small instances whose sizes are equal to or less than 10 pixels. Compared with the previous versions, DOTA-V2.0 (Ding et al., 2022) contains more images collected from Google Earth, GF-2 Satellites, and aerial platforms. In addition, a large number of images are taken under an oblique view and a lower foreground ratio to approach the real-world application scenes. The number of instances has increased to about 1.8 million. Moreover, it contains two test subsets, namely test-dev and test-challenge. The latter comprises a greater number of object instances (around 1.1 million) and more complicated scenes, making the task more challenging.
@@
DIOR-R (Cheng et al., 2022a) is a large-scale dataset that contains 192,518 instances, covering 20 common categories with notable inter-class similarity and intra-class discrepancies. There are 23,463 images chosen carefully from more than 80 countries thereby possessing richer variations in viewpoint, illumination, background, appearance, occlusion, etc. In addition, the GSD ranges from 0.5m to 30m, causing a large range of size variations. Thus, the rich diversity among instances, images, and scales makes this dataset valuable for real-world tasks yet brings about challenges.
@@
Ship Datasets. Recently, a series of ship datasets have drawn wide attention owing to the potential value of ship detection in fishing, maritime security, and national defense. HRSC2016 (Liu et al., 2016b) is one of the most widely used datasets for evaluating algorithms of oriented object detection. It covers more than 25 categories of ships with large varieties of scales, orientations, appearances, shapes, and backgrounds (e.g., sea, port). FGSD (Chen et al., 2020) is a new fine-grained ship detection dataset expanded based on HRSC2016. The instances are classified into 43 categories which are further divided into 4 high-level categories, including submarine, aircraft carrier, civil ship, and warship. Except for ships, a new category named dock is also annotated in this dataset for future research. ShipRSImageNet (Zhang et al., 2021b) is the largest RS dataset for ship detection. It contains 3,435 images collected from xView (Lam et al., 2018), HRSC2016 (Liu et al., 2016b), FGSD (Chen et al., 2020), Airbus Ship Detection Challenge, and Chinese satellites. A total number of 17,573 ships are divided into 50 categories. There are diverse spatial resolutions, scales, aspect ratios, backgrounds, and orientations in this dataset.
7.2 Evaluation Protocol
(1) The predicted label is in accordance with the label of GT object.

### v5

@@ 6.3 Large Aspect Ratio
RS images frequently encompass several categories with extremely large aspect ratios, such as bridges, ships, and harbors. The RIoU between these categories with anchors exhibit significant sensitivity to orientation errors, thereby causing two primary challenges, i.e., spatial misalignment and inaccurate localization.
Accordingly, a series of well-designed assignment strategies and regression loss are proposed to remedy corresponding issues. [208] introduce length-independent IoU (LiIoU). LiIoU intercepts part of the object box along its long side based on the length of the anchor, and subsequently calculates the IoU between the intercepted object box and anchor. It facilitates the assignment of more positive samples to long objects compared to the conventional IoU, thereby enhancing the recall rate. [124] emphasize that the sampling locations of positive samples should be close to the center distribution of the oriented GT. This is because that two horizontal anchors, despite having the same IoU with the external HBB of an oriented GT with large aspect ratio, can differ considerably in their coverage of the GT. Consequently, [124] and [170] both combine the horizontal IoU and sample feature alignment overlap to evaluate the quality of anchor, in which the sample feature alignment overlap is defined as the ratio of the intersection between the horizontal anchor and the oriented GT to the GT itself, reflecting the overlap proportion of oriented object features.
Furthermore, inspired by ATSS [206], SASM [56] and CFL [144] use a monotonic decreasing function of aspect ratio as a weight for the IoU threshold that controls the sample assignment. This approach allows long objects to be assigned a lower IoU threshold. Additionally, several methods [105, 123, 42, 173] construct the weighted orientation loss that depend on aspect ratio, effectively mitigating the effect of aspect ratio on orientation regression.
6.4 Lack of Oriented-Annotated Samples
Relying on massive and laborious annotations, oriented object detection have shown significant advancement in recent years. However, OBB annotation is a time-consuming and expensive process, resulting in many detection datasets that use HBB annotations not providing OBB annotations, thereby restricting the potential scope of the application. To alleviate the annotation burden, researchers have explored two avenues: weakly supervised learning, in which OBB annotations are replaced by HBB-level or even point-level annotations, and semi-supervised learning, in which only a few images from the whole training dataset are annotated.
Mainstream semi-supervised oriented object detection methods commonly follow the pseudo-labeling framework, consisting of a teacher model and a student model. The teacher model, an Exponential Moving Average (EMA) [147] of the student model at historical training iterations, generates pseudo-labels on unannotated images. They are trained iteratively by the following steps: Teacher model provides pseudo-labels for unannotated images in a batch, while student model makes predictions for both annotated and unannotated ones. Then, computing loss for the student model’s predictions. However, the unsupervised nature of the teacher model introduces noise that can mislead the student model’s training, especially considering the arbitrary orientations of objects, which further impacts pseudo-label quality. Therefore, current semi-supervised oriented object detection methods [52, 164, 162] are committed to generating high-quality pseudo-labels.
SOOD [52] introduces two loss functions, termed rotation-aware adaptive weighting (RAW) loss and global consistency (GC) loss. RAW focus on the orientation consistency between each pseudo-label-prediction pair, dynamically weighting each them by their orientation gap. GC measures the global similarity between the pseudo-labels and the predictions, effectively mitigating the noise disturbance and implicitly regularizes object relations. [164] provides an in-depth analysis of the limitation of conventional pseudo-labels and dense pseudo-labels [209] methods. The former adopts a fixed threshold, while the latter uses a fixed quantity, both failing to adaptively select high-quality pseudo-labels. To address this, Global Focal Learning is proposed to judge important regions based on the difference between the prediction of teacher model and student model, guiding the networks to focus more on inconsistent regions during the training. Besides, Pseudo-Siamese Teacher [162] adopts two teacher models updated by different optimizations to improve reliability of pseudo-labels, using Jensen-Shannon divergence to eliminate inconsistent pseudo-labels.
Mainstream weakly-supervised oriented object detection methods commonly consist of multiple branches respectively fed with multiple augmented views of the input image. Then, various consistent losses are designed to align the features or predictions of different views. H2RBox [201], the first HBB annotation-based weakly-supervised method, follows a weakly- and self-supervised angle learning paradigm. The weakly-supervised part calculates regression loss between the external HBB of predicted OBB and horizontal GT, while the self-supervised part measures the consistency of the predicted angles in two views with different rotation augmentation. Based on H2RBox, H2RBox-v2 [195] leverages reflection symmetry to learn the orientation of objects in a self-supervised manner. Furthermore, a CircumIoU loss is designed, empowering H2RBox-v2 compatible with random rotation augmentation.
Compared to OBB and HBB annotations, point annotations show lower costs and higher efficiency333According to https://cloud.google.com/ai-platform/data-labeling/pricing, the cost of point annotations is about lower than HBB and lower than OBB, and their time-consuming is just 1.2x more than image-level annotations. The main challenge of point annotation-based method lies in enabling the model to perceive the orientation and scale of the objects based on point annotations. PointOBB [102] design a resized view (by random scaling) and a rot/flp view (by random rotating or vertical flipping) based on original view. Upon these three views, a scale augmentation module and an angle acquisition module are constructed. The former aims to perceive object scale by improving the consistency of predicted scores between original and resized views, while the latter incorporates self-supervised angle learning to predict angles.
Furthermore, [196] present Point2RBox, including Synthetic Pattern Knowledge Combination and Transform Self-Supervision. The former first generates synthetic patterns with known boxes by sampling around each labeled point, and then overlaid these patterns on the original image, providing the knowledge for network to estimate the size and angle. The latter is similarly to PointOBB, which uses original and transformed (randomly selected from rotate, flip, and scale) view to perceive the size and orientation between objects. In addition, [215] propose a progressive method, named Point-to-Mask-to-HBB-to-OBB (PMHO), to achieve oriented object detection. However, this framework is time-consuming, with each component being optimized independently, relying heavily on the capabilities of well-trained models such as SAM [69].
@@ 6.5 Discussion
Complex background, scale variations, large aspect ratio, and lack of oriented-annotated samples are crucial issues in RS object detection tasks, which evolve more severely when it comes to oriented object detection tasks. As stated above, a number of works have been proposed to tackle these issues from various perspectives, e.g., data augmentation, assignment strategies, re-weighted orientation loss, attention mechanism, self-supervised loss, and pseudo-labeling framework. Unfortunately, exploration into solutions for these issues is far from mature and so further research may be beneficial. For instance, there is still a significant performance gap in detecting small/long and normal objects even for state-of-the-art detectors. On the other hand, the general split-and-detect scheme is inefficient during inference due to too many empty patches that only contain background. Several prior works have provided preliminary consideration on these points, e.g., super-resolution-based object detection [142, 76, 214], focus-and-detect schemes [28, 68]. Additionally, the accuracy of semi/weakly-supervised oriented object detection is still far from satisfactory, lagging significantly behind the fully supervised methods. Combining weakly-supervised and semi-supervised methods may bring new breakthroughs [163].
7 Evaluation Protocol and Datasets
7.1 Evaluation Protocol
(1) The predicted label is in accordance with the label of GT object.
@@ 7.2 Datasets
Recently, several research groups have released dozens of high-quality RS image datasets, each of which dramatically boosts the development of RS object detection. Datasets annotated only with HBBs are not covered here, including DIOR [99], LEVIR [217], NWPU VHR-10 [9], RSOD [180, 81], xView [84], and HRRSD [225]. In addition, several oriented object detection datasets with horizontal views, e.g., text detection datasets [67], and datasets from different modalities, like SAR datasets [167, 87], exhibit significant differences in viewing angles, scenes, and imaging characteristics when compared to optical remote sensing datasets. Thus, this paper will not introduce these datasets. In this subsection, we only focus on introducing optical RS datasets annotated with OBBs, including SZTAKI-INRIA [1], 3K vehicle [93], UCAS-AOD [204], VEDAI [132], HRSC2016 [100], DOTA [172, 30], ShipRSImageNet [233], DIOR-R [19], DroneVehicle [136], FAIR1M [143], and GLH-Bridge [92]. Tab. 3 statistics the parameters of the above RS oriented object detection datasets for intuitive comparison. Given that the emergence of DOTA greatly promotes the development of oriented object detection, we divide the dataset into two parts: early and modern datasets, based on the timeline of DOTA’s introduction. Only the most typical among the above datasets are described in detail due to the space restriction. For more details please refer to Sec. C of Appendix.
Dataset | Publication | Category | Quantity | Instance | GSD | Resolution | |
Early | SZTAKI-INRIA [1] | TPAMI 2012 | 1 | 9 | 665 | - | |
3K vehicle [93] | GRSL 2015 | 1 | 20 | 14,235 | 0.13m | ||
UCAS-AOD [204] | ICIP 2015 | 2 | 2,420 | 14,596 | - | ||
VEDAI [132] | JVCIR 2016 | 9 | 1,210 | 3,640 | 0.125m | ||
HRSC2016 [100] | GRSL 2016 | 25 | 1,070 | 2,976 | 0.42m | ||
Modern | DOTA-V1.0 [172] | CVPR 2018 | 15 | 2,806 | 188,282 | 0.14.5m | |
DOTA-V1.5 | — | 16 | 2,806 | 403,318 | 0.14.5m | ||
DOTA-V2.0 [30] | TPAMI 2022 | 18 | 11,268 | 1,793,658 | 0.14.5m | ||
FGSD [20] | arxiv 2020 | 43 | 5,634 | 2,612 | 0.121.93m | ||
ShipRSImageNet [233] | JSTAR 2021 | 50 | 3,435 | 17,573 | 0.126m | ||
DIOR-R [19] | TGRS 2022 | 20 | 23,463 | 192,518 | 0.530m | ||
DroneVehicle [136] | TCSVT 2022 | 5 | 56,878 | 953,087 | - | ||
FAIR1M [143] | ISPRS 2022 | 37 | 42,796 | 1,000,000 | 0.30.8m | ||
GLH-Bridge [92] | TPAMI2024 | 1 | 6,000 | 59,737 | 0.31.0m |
@@
DOTA [172, 30] contains large quantities of objects with a considerable variety of orientations, scales, and appearances. The images are selected from different sensors and platforms, including Google Earth, GF-2 Satellite, and UAVs. There are three versions of this dataset. The number of images and instances in three versions of DOTA are summarized in Tab. 4. DOTA-V1.0 [172] and DOTA-V1.5 share the same images, which are split into training, validation, and test subsets. As an extension of DOTA-V1.0, DOTA-V1.5 annotates extremely small instances whose sizes are equal to or less than 10 pixels. Compared with the previous versions, DOTA-V2.0 [30] contains more images collected from Google Earth, GF-2 Satellites, and aerial platforms. In addition, a large number of images are taken under an oblique view and a lower foreground ratio to approach the real-world application scenes. The number of instances has increased to about 1.8 million. Moreover, it contains two test subsets, namely test-dev and test-challenge. The latter comprises a greater number of object instances (around 1.1 million) and more complicated scenes, making the task more challenging.
@@
DIOR-R [19] is a large-scale dataset that contains 192,518 instances, covering 20 common categories with notable inter-class similarity and intra-class discrepancies. The previous version of DIOR-R, i.e., DIOR [99], was initially released in 2019 using HBB annotations. Later, in 2021, OBB annotations were added to form the DIOR-R dataset. There are 23,463 images chosen carefully from more than 80 countries thereby possessing richer variations in viewpoint, illumination, background, appearance, occlusion, etc. In particular, it contains some traffic infrastructures due to their significant value in transportation analysis, such as train stations, expressway service areas, and airports, as well as some common categories in the suburbs, such as dam and wind mill. In addition, the GSD ranges from 0.5m to 30m, causing a large range of size variations. Thus, the rich diversity among instances, images, and scales makes this dataset valuable for real-world tasks yet brings about challenges.
@@
Ship Datasets. Recently, a series of ship datasets have drawn wide attention owing to the potential value of ship detection in fishing and maritime security. HRSC2016 [100] is one of the most widely used datasets for evaluating algorithms of oriented object detection. It covers more than 25 categories of ships with large varieties of scales, orientations, appearances, shapes, and backgrounds (e.g., sea, port). FGSD [20] is a new fine-grained ship detection dataset expanded based on HRSC2016. The instances are classified into 43 categories which are further divided into 4 high-level categories, including submarine, aircraft carrier, civil ship, and warship. Except for ships, a new category named dock is also annotated in this dataset for future research. ShipRSImageNet [233] is the largest RS dataset for ship detection. It contains 3,435 images collected from xView [84], HRSC2016 [100], FGSD [20], Airbus Ship Detection Challenge, and Chinese satellites. A total number of 17,573 ships are divided into 50 categories. There are diverse spatial resolutions, scales, aspect ratios, backgrounds, and orientations in this dataset.
DroneVehicle [136] is a large-scale RGB-infrared cross-modal vehicle detection dataset captured by UAVs. This dataset is released to address vehicle detection challenges in smart city traffic management and disaster rescue, especially under conditions of insufficient lighting. The dataset includes two modalities: RGB images and infrared images, with an equal number of images in each modality, collectively forming image pairs. This dual-modal design can provide complementary information under different lighting conditions, e.g., RGB images provide rich color information, while infrared images excel in low-light conditions, unaffected by darkness. Besides, it covers a wide range of scenarios from day to night, including urban roads, rural areas, residential areas, parking lots, etc., ensuring the diversity and practicality of the data.
FAIR1M [143] is currently the largest fine-grained object detection dataset for high-Resolution remote sensing images, containing more than one million instances and over 40,000 images. All instances in this dataset are carefully annotated with OBBs, covering 5 main categories and 37 fine-grained subcategories, such as different types of aircraft, ships, court, road, and vehicles. The images are sourced from different sensors and platforms, with target scenes covering hundreds of typical cities and towns as well as commonly used airports and ports globally, providing rich geographic information and practical application scenarios. The fine-grained annotations, intra-class variations and inter-class variations similarities, large ranges of sizes and orientations, and complex scenes make the dataset extremely challenging, while also promoting the development of object detection in the field of remote sensing.
GLH-Bridge [92] is a large-scale bridge detection dataset comprising 6,000 very-high-resolution RS images sampled from diverse geographic locations around the globe. This dataset covers a wide range of scenarios and bridge types, enhancing its generalizability to real-world situations. Additionally, the various object scales and extreme aspect ratios poses a formidable challenge for oriented object detection methods.
7.3 Discussion
Early datasets often feature a limited number of instances and images, encompassing a narrow range of scenarios. Consequently, detection methods approach saturation in performance on these datasets, making them unable to provide reliable evaluations. Modern datasets generally encompass more challenging and general scenarios with relatively complex backgrounds. They not only cover an astonishing number of instances (up to millions) and fine-grained categories (e.g., FGSD and FAIR1M), but also possess large-scale images (up to pixels), making them highly aligned with real-world application scenarios.
Both early and modern datasets play a pivotal role in propelling oriented object detection methods to new heights, making significant contributions to the field. HRSC2016 [100] is a early benchmark, but as detection methods reached their performance plateau on this dataset, it gradually fell out of favor. Although subsequent datasets, e.g., FGSD [20] and ShipRSImageNet [233], attempt to expand upon it, they failed to garner sufficient attention due to its relatively limited number of instances compared to those larger-scale datasets. DOTA-V1.0 [172], as the most representative datasets for oriented object detection, serve as the most commonly used benchmark for evaluating the performance of detection methods. Subsequently, DOTA-V2.0 [30], characterized by its large scale and high level of challenge, and FAIR1M [143], which focuses on fine-grained oriented object detection, gradually become new benchmarks for evaluating the performance of cutting-edge methods.
Modern large-scale datasets provide a solid data foundation for the deployment and implementation of real-world applications. By utilizing large-scale datasets for pre-training and transfer learning, the development costs and time can be significantly reduced, while the recognition accuracy of oriented object detection models can be improved, facilitating their application in various fields.
City management. With the acceleration of urbanization and the continuous expansion of urban areas, traditional ground traffic monitoring systems gradually reveal their limitations in terms of delayed response. Leveraging their remarkable flexibility, UAVs are widely applied in traffic dispersion and traffic flow monitoring [136, 161], emerging as a cutting-edge force in urban traffic management.
Industrial inspection. Industrial facilities such as bridges and wind turbines often require a significant amount of labor and time for inspections, posing challenges for manual inspections [19, 92]. Incorporating UAVs or satellites and intelligent detection technology can significantly improve inspection efficiency and ensure personnel safety.
Port management. Utilizing advanced image processing algorithms and object detection technology, various objects within the port area, such as ships and port facilities, can be automatically identified from satellite imagery, thereby optimizing port operational efficiency and enhancing safety [100, 233].
Security surveillance. The detection of critical objects in satellite imagery, such as airplane and airports, plays a vital role in security surveillance [30].